OpenStudy (anonymous):

What is the value of r + s + t when r = 26, s = –12, and t = –8? A. 6 B. 22 C. 30 D. 46

OpenStudy (anonymous):

A.

OpenStudy (zale101):

r + s + t when r = 26, s = –12, and t = –8 the question is telling you to replace r by 26, s by -12, and t by -8 26+(-12)+(-8)=? 26-12-8=?